OpenStudy (anonymous):

You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly? Type your answer below using complete sentences.

OpenStudy (campbell_st):

this is a binomial probability question let p be a correct answer with a probability of 1/4 and q be an incorrect answer with a probability of 3/4 the probability of 3 correct from 5 can be found using \[^nC_{k} \times p^k \times q^{n - k}\] I have already defined p and q you need to know n = 5, k = 3 and C means combination... hope it helps

OpenStudy (anonymous):

I still am lost can you explain it a little more clearly

OpenStudy (campbell_st):

ok... so here it is with the appropriate substitutions and binomial probabilities uses the probability of success (p) 1/4 (providing th correct answer from 4 choices) and probability of failure (q) 3/4 ( 3 incorrect answers out of 4) so the answer is \[^5C_{3} \times (\frac{1}{4})^3 \times (\frac{3}{4})^{5 - 3}\] just calculate it out...

OpenStudy (anonymous):

so what would C be

OpenStudy (campbell_st):

C stands for combination.... \[^5C_{3} = \frac{5!}{3! \times (5 - 3)!}\]

OpenStudy (anonymous):

Its still doesnt make any sense

OpenStudy (campbell_st):

ok... here have you studied binomial probabilities.... combinations and permuatations...?

OpenStudy (anonymous):

yes accept the last one

OpenStudy (campbell_st):

ok.... well this question is asking you to use binomial probability to calculate the solution you may have seen the formula as \[\left(\begin{matrix}n \\ k\end{matrix}\right) \times (p)^k \times (q)^{n - k}\] in binomial probability there are only 2 things that can happen getting a right answer....P(right) = 1/4 and getting a wrong answer ....P(wrong) =3/4 you need to use both of these pieces of information to find the probability of 3 correct from 5... is that any clearer...

OpenStudy (anonymous):

nope

OpenStudy (campbell_st):

ok... here are some notes and examples I just googled which goes through Binomial probabilities. hope it helps... then comeback and look at the question and solution

OpenStudy (campbell_st):

oops forgot the link http://www.regentsprep.org/Regents/math/algtrig/ATS7/BLesson.htm

OpenStudy (anonymous):

wow doesnt help a bit

OpenStudy (campbell_st):

ok.... well I think you need to sit down with your instructor and go though it... I'm confident my solution is correct...

OpenStudy (anonymous):

ok so what would you think the answer would be?